The current price of Estelle Corporation stock is $ 28.00 In each of the next two​ years, this stock price will either go up by 21% or go down by 21%. The stock pays no dividends. The​ one-year risk-free interest rate is 7.6% and will remain constant. Using the Binomial​ Model, calculate the price of a​ one-year put option on Estelle stock with a strike price of $28.00.
Answer #1
High price=Current price*up move=28*1.21=33.88
Low price=Current price*down move=28*0.79=22.12
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.076*1)-0.79)/(1.21-0.79)=0.68801
Put option payoff at high price (payoff H)
=Max(Strike price-High price,0)
=Max(28-33.88,0)
=Max(-5.88,0)
=0
Put option payoff at low price (Payoff L)
=Max(Strike price-low price,0)
=Max(28-22.12,0)
=Max(5.88,0)
=5.88
Price of Put option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.076*1)*(0.688006*0+(1-0.688006)*5.88)
=1.7